If a + 2b is equal to 125 percent of 4b, what is the value of a/b?
Answer by ankor@dixie-net.com(22740)

"a + 2b is equal to 125 percent of 4b,"
:
a + 2b = 1.25(4b)
a + 2b = 5b
a = 5b - 2b
a = 3b
:
what is the value of a/b?
Substitute 3b for a and you have:
(3b)/b = 3;
:
(b = 1 and a = 3)